OpenStudy (omgthatgirl):

Simplify 2^1/3 times 4^1/3 using the radical form

OpenStudy (omgthatgirl):

A. 512 B. 2^2/3 C. 2. D. 8

OpenStudy (bibby):

rewrite the 2nd one in terms of a common base with the 1st term use the following information: \(\huge (a^b)^c=a^{(b*c)}\\\huge a^b*a^c=a^{(b+c)}\)

OpenStudy (omgthatgirl):

Ok

OpenStudy (omgthatgirl):

It would be 8

OpenStudy (bibby):

I didn't get that, show your work

OpenStudy (omgthatgirl):

Well I multiplied 4 and 2 and got 8 1/3=1/3 so I left it that way

OpenStudy (bibby):

you can't do that. that's why I said you had to convert it. \(\huge (a^b)^c=a^{(b*c)}\\\huge a^b*a^c=a^{(b+c)}\) note that the bases are the same for the 2nd equation. that means we have to rewrite 4 as a po wer of 2

OpenStudy (omgthatgirl):

Ok so the exponents would equal 2/3

OpenStudy (bibby):

not really, you're not listening to what I'm saying. \(\large 4^\frac{1}{3}=2^?\)

OpenStudy (jdoe0001):

\(\Large {\bf 2^{\frac{1}{3}}\cdot 4^{\frac{1}{3}}\qquad {\color{brown}{ 4=2^2}}\qquad 2^{\frac{1}{3}}\cdot (2^2)^{\frac{1}{3}}\implies (2)^{2\cdot \frac{1}{3}} }\) can you see wht bibby is saying now?

OpenStudy (omgthatgirl):

1/3?

OpenStudy (bibby):

lol, sorry, I'm never sure what I should illustrate and what I should be explaining. thanks jode anyhow, if \( 4=2^2\) and \(\large(a^b)^c=a^{(b*c)}\) then \(4^frac{1}{3}=(2^2)^\frac{1}{3}\)

OpenStudy (omgthatgirl):

Thank you both

OpenStudy (bibby):

\(\huge 2^\frac{1}{3} \times 2^\frac{2}{3}=?\)

OpenStudy (omgthatgirl):

2